solve use reduced fractions not decimals
x-2/2+1/3=x/6

Answer by checkley77(12844)

x-2/2+1/3=x/6
X-1+1/3-X/6=0
X-X/6=1-1/3
6X/6-X/6=3/3-1/3
5X/6=2/3 CROSS MULTIPLY
3*5X=6*2
15X=12
X=12/15
X=4/5 ANS.
PROOF:
4/5-1+1/3=(4/5)/6
12/15-15/15+5/15=(4/5)*1/6
2/15=4/30
2/15=2/15